OpenStudy (anonymous):

2/3 (c.18)=7 Help Solve equation

OpenStudy (anonymous):

is c.18 c multiply by 18

OpenStudy (anonymous):

Yes It is

OpenStudy (anonymous):

2/3*18c = 7 2*18c = 21 36c = 21 c = 21/36 c = 7/12

OpenStudy (anonymous):

Is that the answer

OpenStudy (anonymous):

Wit how do you get the 18c

OpenStudy (anonymous):

by multiplying 18 with c

OpenStudy (anonymous):

Oh thank you i thought you couldn't multiply unless they are like terms.